Ben bowled 127 and 220 in his first two games. What must he bowl in his third game to have an average of at least 170?

Answer by checkley79(3341)

(127+220+X)/3>=170 CROSS MULTIPLY.
127+220+X=170*3
347+X>=510
X>=510-347
X>=163 IS NEEDED ON THE THIRD GAME TO AVERAGE AT LEAST 170
PROOF:
(127+220+163)/3>=170
510/3>=170
170>=170